# 3D-Guided Scalable Flow Matching for Generating Volumetric Tissue Spatial Transcriptomics from Serial Histology

### v1

@@ 5.3 HoloTea captures expression morphology of anatomical marker genes
We trained a model for the 211 marker genes on the in-house embryo dataset, see LABEL:sec:imputing-marker-genes.  Tab. 3 visualizes the expression patterns for some marker genes of two different anatomical regions, the basal plate (CNMD and COL2A1) and the cardinal vein (CEMIP and SEMA3E), showing ground truth and imputed expression across 3 of the 26 sections held out during training. It is notable that the imputed expression patterns closely recapitulate the ground truth expression, while reducing expression variance in the low-count regime. In particular, for the basal plate markers, the imputed expression patterns exhibit segmentation along the spinal cord, and imputed expression in the anatomical region of the cardinal vein is observed for its marker genes. Further, while secondary regions with lower expression are also recapitulated in the imputed data, binary shot noise is suppressed.
Jointly clustering the ground truth and imputed expression of all 26 unseen sections (Tabs. 3, LABEL:sec:clustering-marker-genes, LABEL:suppl-fig:clustering-gt, LABEL:suppl-fig:clustering-imp and LABEL:suppl-fig:clustering-summary) shows that imputed expression clusters recapitulate those of the ground truth, with a normalized mutual information of 0.524. The ground truth clusters have a lower neighbor consistency of $0.775\pm 0.109$ than the imputed clusters, $0.880\pm 0.045$, indicating that the imputed clusters are more spatially contiguous.
Overall, these results suggest that imputing expression on interleaved sections is feasible. Additional details about experimental results can be found in LABEL:app:results_details.
@@ 6 Ablation
To quantify the contribution of each architectural component, we compare four variants of our model on the in-house embryo dataset and evaluate them on the even-slice hold-out. We sweep over marker panel sizes of 50, 100, 200, and 500 genes, and report spot-level mean squared error (MSE), mean absolute error (MAE), and Pearson correlation coefficient (PCC) in Table 5. The four regimes are: (i) Vanilla STFlow, our base model; (ii) Vanilla + Learned Prior, which augments STFlow with a learned ZINB prior; (iii) ControlNet + Learned Prior, which additionally uses ControlNet-style conditioning; and (iv) ControlNet + CrossCosine + Learned Prior, which further adds our cross-slide cosine branch.
@@
Adding ControlNet on top of the learned prior yields consistent gains over both Vanilla STFlow and the prior only model. For the 50 gene panel, PCC increases from 0.693 (vanilla) and 0.691 (prior only) to 0.701, while MSE decreases to 0.154, the best among all regimes at this panel size. Similar trends hold at 100 and 200 genes, where ControlNet improves PCC to roughly 0.690.69-$0.70$ with reduced MSE/MAE. Even for the widest 500 gene panel, ControlNet remains beneficial. Qualitative comparison of predicted spatial expression patterns for gene AHSP, for our ControlNet-based model (top row) and STFlow (bottom row) across flow steps (1, 3, 5), with ground-truth expression shown on the right Fig. 3. Our model, which is equipped with ControlNet and cross-slide conditioning, more faithfully recovers the high-expression clusters and fine-grained structures than STFlow, visually reinforcing the quantitative gains reported in the ablation study (Table 5).
@@
Table 5: Spot-level performance on the even-slice hold-out split for different ablation settings on the in-house embryo dataset. For each panel size, the best value is shown in bold. Lower is better for MSE/MAE; higher is better for PCC. Due to space constraints, the Learned Prior row is omitted; the full table is provided in the Supplementary Material LABEL:app:results_details.

### v2

@@ 5.3 HoloTea captures expression morphology of anatomical marker genes
We trained a model for the 211 marker genes on the in-house embryo dataset, see Tab. S10.  Tab. 3 visualizes the expression patterns for some marker genes of two different anatomical regions, the basal plate (CNMD and COL2A1) and the cardinal vein (CEMIP and SEMA3E), showing ground truth and imputed expression across 3 of the 26 sections held out during training. It is notable that the imputed expression patterns closely recapitulate the ground truth expression, while reducing expression variance in the low-count regime. In particular, for the basal plate markers, the imputed expression patterns exhibit segmentation along the spinal cord, and imputed expression in the anatomical region of the cardinal vein is observed for its marker genes. Further, while secondary regions with lower expression are also recapitulated in the imputed data, binary shot noise is suppressed.
Jointly clustering the ground truth and imputed expression of all 26 unseen sections (Tabs. 3, S10, S2, S3 and S3) shows that imputed expression clusters recapitulate those of the ground truth, with a normalized mutual information of 0.524. The ground truth clusters have a lower neighbor consistency of $0.775\pm 0.109$ than the imputed clusters, $0.880\pm 0.045$, indicating that the imputed clusters are more spatially contiguous.
Overall, these results suggest that imputing expression on interleaved sections is feasible. Additional details about experimental results can be found in Sec. S3.
@@ 6 Ablation
To quantify the contribution of each architectural component, we compare four variants of our model on the in-house embryo dataset and evaluate them on the even-slice hold-out. We sweep over marker panel sizes of 50, 100, 200, and 500 genes, and report spot-level mean squared error (MSE), mean absolute error (MAE), and Pearson correlation coefficient (PCC) in Table S10. The four regimes are: (i) Vanilla STFlow, our base model; (ii) Vanilla + Learned Prior, which augments STFlow with a learned ZINB prior; (iii) ControlNet + Learned Prior, which additionally uses ControlNet-style conditioning; and (iv) ControlNet + CrossCosine + Learned Prior, which further adds our cross-slide cosine branch.
@@
Adding ControlNet on top of the learned prior yields consistent gains over both Vanilla STFlow and the prior only model. For the 50 gene panel, PCC increases from 0.693 (vanilla) and 0.691 (prior only) to 0.701, while MSE decreases to 0.154, the best among all regimes at this panel size. Similar trends hold at 100 and 200 genes, where ControlNet improves PCC to roughly 0.690.69-$0.70$ with reduced MSE/MAE. Even for the widest 500 gene panel, ControlNet remains beneficial. Qualitative comparison of predicted spatial expression patterns for gene AHSP, for our ControlNet-based model (top row) and STFlow (bottom row) across flow steps (1, 3, 5), with ground-truth expression shown on the right Fig. 3. Our model, which is equipped with ControlNet and cross-slide conditioning, more faithfully recovers the high-expression clusters and fine-grained structures than STFlow, visually reinforcing the quantitative gains reported in the ablation study (Table S10).
@@
Table 5: Spot-level performance on the even-slice hold-out split for different ablation settings on the in-house embryo dataset. For each panel size, the best value is shown in bold. Lower is better for MSE/MAE; higher is better for PCC. Due to space constraints, the Learned Prior row is omitted; the full table is provided in the Supplementary Material S3.
@@ 7 Conclusion
Supplementary Material
Table S1: Performance across the HER2, ST-Data, and in-house embryo datasets on highly variable gene sets.
HER2 ST-Data (images) In-house embryo
Model MSE (↓) MAE (↓) PCC (↑) MSE (↓) MAE (↓) PCC (↑) MSE (↓) MAE (↓) PCC (↑)
SPOT GENE SPOT GENE SPOT GENE
UNI $1.093$ $0.859$ $0.509(94)$ $0.259(231)$ $0.769$ $0.699$ $0.569(77)$ $0.332(149)$ $0.261$ $0.296$ $0.506(224)$ $0.613(203)$
CIGA $0.865$ $0.738$ $0.494(135)$ $0.465(143)$ $0.7825$ $0.706$ $0.541(83)$ $0.247(137)$ $0.449$ $0.438$ $0.399(194)$ $0.427(229)$
Gigapath $0.611$ $0.615$ $0.581(126)$ $0.576(161)$ $0.674$ $0.649$ $0.597(76)$ $0.364(145)$ $0.314$ $0.322$ $0.457(223)$ $0.554(225)$
[…]
S1 Dataset Details
S1.1 In-house embryo dataset.
[…]
The spatial transcriptomic data were processed using the Space Ranger v2.10 pipeline with standard quality control procedures. In total, the dataset comprises 327,532 spatial spots, with a median of 7,618 spots under tissue. The average median number of genes detected per spot across all sections is approximately 7,173, providing high spatial resolution across the developing embryo.
S1.2 Public datasets.
HER2 dataset.
For the HER2 cohort we used the spatial transcriptome sequencing data from eight HER2 positive human breast tumors, where each tumor contributes three or six spatial transcriptomics sections, giving a total of 36 profiled sections. These samples were generated with the first generation Spatial Transcriptomics in situ capturing arrays and sequenced on an Illumina NextSeq 550 platform, yielding whole transcriptome counts for each capture spot. The original study combines these data with single cell RNA sequencing from additional HER2 positive tumors to deconvolve each spot into proportions of major cell types, including epithelial, myeloid, T, B, plasma, endothelial and cancer associated fibroblast populations, and reports that a typical spot aggregates roughly 0 to 200 cells. A pathologist manually annotated one section per tumor into regions such as in situ cancer, invasive cancer, immune infiltrate, adipose tissue and connective tissue, and expression based clustering plus pathway enrichment were then used to define cancer and immune core signatures and to map tertiary lymphoid like structures and type I interferon related immune niches across all 36 sections.
ST-data (Integrating spatial gene expression and breast tumour morphology via deep learning).
The ST-data cohort originates from the work of He et al., who introduced ST Net using a spatial transcriptomics dataset that links histopathology to local gene expression in breast cancer. The public release contains triplicate spatial transcriptomics sections from 23 patients spanning luminal A, luminal B, HER2-positive, and triple negative subtypes, for a total of 68 sections and 30,612 spatial locations with matched hematoxylin and eosin whole slide images. Each capture spot has a diameter of about 100 micrometer on the original in situ capturing array, and the corresponding histology was digitized at 20x, so that a 224 by 224 pixel image patch centered on each spot covers roughly 150 by 150 micrometer of tissue. The sequencing depth in the original study results in a median of around 3,400 counts and several thousand detected genes per spot, and ST Net was trained to predict the log expression of a curated panel of 250 target genes from these image patches using a leave one patient out cross validation protocol. This dataset therefore, provides a well-curated benchmark with consistent pre-processing, subtype labels, and tumor or normal tissue annotations, which we reuse as a standard testbed for evaluating our method on two-dimensional sections.
S1.3 Spatial alignment of Visium slides
Data alignment.
For the public datasets (HER2 and ST-data), we obtained the raw data from the original studies’ official repositories and did not perform any additional alignment beyond the preprocessing carried out by the original authors. In contrast, for our in-house embryo dataset, we applied the two-stage alignment workflow described below.
Stage 1: Pre-alignment with MOSCOT.
[…]
Stage 2: Fine registration with pystackreg.
The pre-aligned image stack served as input for fine-grained image registration implemented with the Python package pystackreg. We used the green (G) channel of the H&E images providing the highest contrast for alignment. Affine registration (translation, rotation, scaling, shear) was performed sequentially: each section was registered to its immediate predecessor (e.g., section 2 to section 1, section 3 to section 2, and so on). This produced a finely registered image stack and an updated set of transformation matrices.
Final spot-coordinate update.
To obtain Visium spot coordinates corresponding to the fully registered images, we applied the transformation matrices from the fine-registration step to the MOSCOT-aligned spot positions, yielding the final coordinates used in downstream analyses. The implementation will be released following peer review.
S1.4 Gene panel definitions.
For completeness, we summarize the three gene panels used throughout the main paper.
  • HVG panel. A highly variable gene (HVG) panel obtained per dataset using a standard Scanpy pipeline (library-size normalization, $\log(1+x)$ transform, and variance ranking). In the main paper we report results for the HVG-250 configuration.
  • Custom panel. A dataset-specific panel constructed by combining spatially variable genes and genes of prior biological interest (for example, pathway genes and marker sets suggested in the original dataset releases).
  • Marker panel. A curated set of $211$ marker genes associated with $106$ anatomical regions in the embryo stack. This panel is used for the clustering and neighbor-consistency analysis described in the main paper.
These panels correspond to the sets referred to as (i) HVG, (ii) custom, and (iii) marker genes in the main text.
S2 Experiments Details
Spot-centric visual features. For each spatial spot, we extract a $224{\times}224$ image patch centered at the spot from the whole-slide image, pass it through a frozen pre-trained foundation encoder, and use the resulting global embedding as the node feature for that spot.
The key hyperparameters used to train the model are summarized in the table S2.
Dataset splits. For all public datasets, we follow the official train/test partitions used in the ASIGN paper [47] to ensure fair comparisons. Unless otherwise noted, all metrics are reported on the held-out sections.
Baseline training protocol.
For all baselines built on pathology foundation encoders (UNI, CIGA, and GigaPath) we follow a consistent training protocol that mirrors HoloTea as closely as possible. The pathology encoder weights are kept frozen, and we train only the downstream ZINB decoder and, where applicable, the flow-matching components, using the optimization hyperparameters listed in Table S3. This ensures that differences in performance primarily reflect the choice of encoder and our 3D-aware architecture rather than differences in training budget. For the ASIGN (3D) baseline we adapt the official configuration to our preprocessing and normalization pipeline and train until validation performance plateaus or GPU memory is exhausted on the largest embryo stack.
Alternative pathology encoders. In addition to the UNI pathology backbone in the main experiments, we also study the effect of swapping the image encoder while keeping the rest of the pipeline fixed. Specifically, in the appendix we repeat our experiments with two frozen H&E encoders: (i) the self-supervised contrastive histopathology model of Ciga et al. [4], pretrained on large collections of unlabeled whole-slide images; and (ii) the GigaPath whole-slide foundation model for digital pathology [43]. For both variants, we extract patch features with the pretrained encoder, freeze all its weights, and train only the ZINB decoder (and, where applicable, the flow-matching components) using the same optimizer, learning rate schedule, and training budget as in the UNI setup. This yields image-only encoder–decoder baselines that isolate the impact of the pathology backbone on 3D imputation performance.
Global context tokens $m$.
All experiments use the global context mechanism described in the main paper, parameterized by the number of inducing tokens $m$. For each dataset we perform a small sweep from
$m\in\{0,4,8,16,32,64,128,256\}$
on a held-out split (see Tables S6–S8), and then select $m$ per dataset based on performance and the trade-off between accuracy and compute. In practice, moderate values of $m$ already capture most of the benefit, while very large values provide only marginal additional gains at a higher computational cost.
Table S2: Core model hyperparameters (defaults).
Parameter Default
Backbone layers $L$ 4
Hidden dimension $d$ 128
Pairwise / edge dimension 128
Attention heads 4
$k$ nearest neighbors 8
Dropout / Attention dropout 0.2 / 0.2
Global context (inducing tokens $m$) based on hyperparameter search
Table S3: Training and inference hyperparameters.
Parameter Default / Setting
Batch size (slide-level) 2
Epochs 100
Optimizer / Learning rate Adam / $5\times 10^{-4}$
Gradient clipping (L2) 1.0
Time conditioning Sinusoidal $\rightarrow$ MLP
Inference steps ($S$) 5
Table S4: Priors and optional modules.
Component Default / Setting
Prior sampler ZINB (default)
ZINB parameters total_count =1=1, logits =0.1=0.1, zi_logits =0.0=0.0
Learned ZINB prior ($\phi_{e}$) optional (frozen at train time)
Spatial–empirical prior optional (k=128k{=}128 neighbors across $z\!\pm\!1$)
ControlNet conditioning optional; channels =32=32, grid =64=64, scale =1.0=1.0
Table S5: Foundation encoders and node-feature dimensions. Encoders are frozen.
Encoder Patch size Output dim
UNI v1 (official) $224{\times}224$ 1024
GigaPath $224{\times}224$ 1536
CIGA $224{\times}224$ 512
S3 Results Details
Complexity and dataset scale.
The embryo dataset used in our experiments consists of 327,532327{,}532 spatial spots across the 3D stack (see Table S9). The complexity numbers reported in Table 4 of the main paper are measured per slide within this stack and should be interpreted in the context of this overall scale: the global attention backbone with $m$ inducing tokens enables training and inference to scale approximately linearly with the number of spots per slide and with the depth of the stack.
Randomness and variability.
Unless otherwise stated, all metrics in the main paper are reported for a single trained model per configuration. For each metric we aggregate over all spots or all genes in the corresponding test set, which substantially reduces statistical noise. Additional runs with different random seeds on a subset of configurations showed only minor fluctuations in the reported metrics, so we omit per-seed averages for brevity and focus on the main configuration in the tables.
Clustering
Clustering of true and predicted counts was performed using routines implemented in the scanpy package. First, counts were normalized to 10 00010\,000 counts per cell, and then log1p-transformed. The transformed counts were projected to the top-50 principal components and 15-nearest neighbors were determined. Finally, clustering analysis was performed using the Leiden community detection algorithm with a resolution parameter of 1.
Neighbor consistency
Neighbor consistency $NC$ for the spot-neighbor graph was calculated according to
$NC=\sum_{(v,w)\in E}\frac{\mathbbm{1}_{c(v)=c(w)}}{\left|E\right|},$ (19)
where $E$ denotes the set of neighboring vertex pairs and $c(v)$ indicates the cluster assignment of vertex $v$.
Effect of $m$ (EMBRYO).
[…]
Interestingly, small values of $m$ already capture most of the gain. At m=4m{=}4, we obtain MAE $0.237$, MSE $0.207$, PCC SPOT $0.571\pm 0.244$, and PCC GENE $\mathbf{0.681}\pm 0.171$, recovering a large fraction of the improvement over m=0m{=}0. For larger $m\in[32,256]$, we observe diminishing returns: errors continue to decrease slightly (e.g., MSE $0.206\to 0.202$), while correlations remain effectively stable within $\sim\!0.002\!-\!0.005$. Taken together, these results indicate that $m$ controls a useful capacity/aggregation trade-off, but relatively small context sizes (e.g., m=416m{=}4\!-\!16) already offer a strong balance between accuracy and efficiency.
Effect of $m$ (ST).
On the ST dataset (Table S7), increasing $m$ yields clear accuracy gains up to a moderate context. Relative to m=0m{=}0 (MSE $0.443$, MAE $0.532$, PCC SPOT $0.775\pm 0.080$, PCC GENE $0.540\pm 0.142$), performance improves steadily and peaks around m=32m{=}32: MSE reaches $\mathbf{0.382}$ (0.061-0.061, $\sim\!13.8\%$ reduction), MAE $\mathbf{0.479}$ (0.053-0.053, $\sim\!10.0\%$ reduction), while correlations also increase to $\mathbf{0.786}$ (SPOT; +0.011+0.011) and $\mathbf{0.572}$ (GENE; +0.032+0.032). Notably, most of these gains are already realized by m=816m{=}8\!-\!16 (e.g., MSE $0.392$, MAE $0.480$ at m=16m{=}16), after which returns diminish and performance slightly plateaus ($m{\geq}32$).
Effect of $m$ (HER2).
For HER2 , the effect of $m$ is modest. Compared to m=0m{=}0 (MSE $0.477$, MAE $0.534$, PCC SPOT $0.740\pm 0.125$, PCC GENE $0.620\pm 0.146$), the best results occur at m=16m{=}16: MSE $\mathbf{0.462}$ (0.015-0.015, $\sim\!3.1\%$ reduction), MAE $\mathbf{0.530}$ (0.004-0.004, $\sim\!0.7\%$ reduction), PCC SPOT $\mathbf{0.741}$ (+0.001+0.001) and PCC GENE $\mathbf{0.623}$ (+0.003+0.003) Table S8. Beyond $m{\approx}16$, curves remain essentially flat. We speculate this limited sensitivity arises from dataset and number of spot size: as summarized in Table S9, HER2 contains relatively few spots, so enlarging $m$ adds little additional neighborhood context. In regimes with sparse spatial neighborhoods, the proposed module may thus offer less scope for improvement, consistent with the small but nonzero gains observed here.
Table S6: Effect of $m$ (powers of two) on performance for the EMBRYO dataset. Lower is better for MAE/MSE; higher is better for PCC SPOT / PCC GENE. The m=0m{=}0 row is the provided baseline.
$m$ MSE $\downarrow$ MAE $\downarrow$ PCC SPOT (mean $\pm$ std) $\uparrow$ PCC GENE (mean $\pm$ std) $\uparrow$
0 0.255 0.270 0.557 $\pm$ 0.248 0.663 $\pm$ 0.166
4 0.207 0.237 0.571 $\pm$ 0.244 0.681 $\pm$ 0.171
8 0.212 0.241 0.570 $\pm$ 0.246 0.680 $\pm$ 0.173
16 0.210 0.240 0.572 $\pm$ 0.246 0.675 $\pm$ 0.174
32 0.206 0.236 0.567 $\pm$ 0.244 0.680 $\pm$ 0.173
64 0.204 0.236 0.568 $\pm$ 0.245 0.679 $\pm$ 0.170
128 0.203 0.235 0.569 $\pm$ 0.242 0.681 $\pm$ 0.173
256 0.202 0.234 0.569 $\pm$ 0.243 0.679 $\pm$ 0.177
Table S7: Effect of $m$ on performance for the ST dataset. Lower is better for MAE/MSE; higher is better for PCC SPOT / PCC GENE.
$m$ MSE $\downarrow$ MAE $\downarrow$ PCC SPOT (mean $\pm$ std) $\uparrow$ PCC GENE (mean $\pm$ std) $\uparrow$
0 0.443 0.532 0.775 $\pm$ 0.080 0.540 $\pm$ 0.142
4 0.412 0.502 0.781 $\pm$ 0.014 0.561 $\pm$ 0.101
8 0.403 0.491 0.782 $\pm$ 0.012 0.563 $\pm$ 0.098
16 0.392 0.480 0.784 $\pm$ 0.012 0.571 $\pm$ 0.012
32 0.382 0.479 0.786 $\pm$ 0.012 0.572 $\pm$ 0.038
64 0.384 0.481 0.780 $\pm$ 0.045 0.569 $\pm$ 0.019
Table S8: Effect of $m$ on performance for the HER2 dataset. Lower is better for MAE/MSE; higher is better for PCC SPOT / PCC GENE.
$m$ MSE $\downarrow$ MAE $\downarrow$ PCC SPOT (mean $\pm$ std) $\uparrow$ PCC GENE (mean $\pm$ std) $\uparrow$
0 0.477 0.534 0.740 $\pm$ 0.125 0.620 $\pm$ 0.146
4 0.473 0.533 0.739 $\pm$ 0.010 0.618 $\pm$ 0.013
8 0.471 0.533 0.740 $\pm$ 0.011 0.62 $\pm$ 0.037
16 0.462 0.530 0.741 $\pm$ 0.023 0.623 $\pm$ 0.012
32 0.472 0.533 0.740 $\pm$ 0.021 0.621 $\pm$ 0.023
Table S9: Average number of spots per dataset.
HER2 ST EMBRYO
378 482 6353
Table S10: Spot-level performance on the even-slice hold-out split for different ablation settings on the in-house embryo dataset. For each panel size, the best value is shown in bold. Lower is better for MSE/MAE; higher is better for PCC.
Regime Genes MSE $\downarrow$ MAE $\downarrow$ PCC(GENE) $\uparrow$
Vanilla STFlow $50$ $0.160$ $0.107$ $0.693$
$100$ $0.177$ $0.120$ $0.688$
$200$ $0.176$ $0.119$ $0.673$
$500$ $0.164$ 0.124 $0.646$
Vanilla + Learned Prior $50$ $0.181$ $0.127$ $0.691$
$100$ $0.175$ $0.116$ $0.690$
$200$ $0.161$ $0.114$ $0.680$
$500$ $0.164$ $0.134$ $0.642$
ControlNet + Learned Prior $50$ 0.154 $0.096$ 0.701
$100$ $0.169$ $0.117$ $0.692$
$200$ $0.155$ $0.108$ $0.694$
$500$ $0.164$ $0.127$ $0.651$
$100$ 0.158 0.112 0.697
$200$ 0.147 0.108 0.698
$500$ 0.162 $0.128$ 0.654
[…]
[ head to column names, tabular= c@ S[table-format=3.0] S[table-format=1.3, round-mode=places, round-precision=3] S[table-format=1.3, round-mode=places, round-precision=3] , table head= Section GT Imp
, table foot= mean 0.775250 0.880854
std 0.109230 0.044669
[…]
Figure S1: Neighbor consistency of joint clustering across 26 unseen sections. GT: ground truth, Imp: imputed, std: standard deviation.
Figure S2: Joint clustering of the ground truth expression across all 26 sections not seen during training.
Figure S3: Joint clustering of the imputed expression across all 26 sections not seen during training.
Figure S4: Imputed gene panels per datasets HER2, ST-Dataset, In-house Embryo HVG and Custom